OpenStudy (anonymous):

Which is an equation of the given line in standard form? The points are (-4,-1) and ( 1 1/2, 2) a. –11x + 6y = –13 b. –6x +11y = 13 c. –6x + 11y = –13 d. –6x + 7y = –17

OpenStudy (caozeyuan):

Since this is a MCQ, not calculation, you can just plug in and see which one fits

OpenStudy (caozeyuan):

e.g. -11*-4+6*-1=44-6 /= 13, so A is wrong